the probability of bomb hitting the bridge is 1/5...2 bombs are enough to destroy the bridge..if six bombs are fired


what is the probability that the bridge is destroyed?


in evaluating this cant we use bernoully distribution for finding probability of exactly 2 bombs hitting the bridge.


but the answer doesnt match....

Grade:12

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Please find the answer to your question below
Here given n=6 and p=1/5 so q=1-p=4/5
Bridge gets destroyed when 2 or more bombs hit it
sothe probability that the bridge is destroyed is
=P(2)+P(3)+P(4)+P(5)+P(6)=1-[P(0)+P(1)]
=1-[6c0(1/5)0(4/5)6+6c1(1/5)(4/5)5]
=0.345